water is flowing in a river at 2.5 m÷s.the river is 40m wide with an average depth of 6.0m.compute the power available from the water current.

Grade:12

3 Answers

Arpit Kumar
10 years ago
Take the time of flow to be 1 second. Volume of water=2.5*40*6=600m^3 Density of water is 1000 kg/m^3 Mass of water=600000 kg Kinetic Energy of water=1/2*600000*2.5*2.5=1875000 J Power generated=1875000 W
lokesh palingi
10 years ago
THANK U SIR,
chaitnyakishore
10 years ago
Power(P)=work/time =(potential energy+kinetic energy)/timet =[mgh+(mv²/2)]/t =[m/t]*[gh+(v²/2)] Therefore P =[density(d)*volume(V)/t]*[gh+(v²/2)] But Volume/time=area(A)*velocity(v) Therefore P=[dAv]*[gh+(v²/2)] =[1000*40*6*2.5]*[g(0)+{(2.5)²/2}] =1000*240*2.5*3.125 =1875000W =1.875MegaWatts.